### v1

# Graph of AI Ideas: Leveraging Knowledge Graphs and LLMs for AI Research Idea Generation

Abstract
Reading relevant scientific papers and analyzing research development trends is a critical step in generating new scientific ideas. However, the rapid increase in the volume of research literature and the complex citation relationships make it difficult for researchers to quickly analyze and derive meaningful research trends. The development of large language models (LLMs) has provided a novel approach for automatically summarizing papers and generating innovative research ideas. However, existing paper-based idea generation methods either simply input papers into LLMs via prompts or form logical chains of creative development based on citation relationships, without fully exploiting the semantic information embedded in these citations. Inspired by knowledge graphs and human cognitive processes, we propose a framework called the Graph of AI Ideas (GoAI) for the AI research field, which is dominated by open-access papers. This framework organizes relevant literature into entities within a knowledge graph and summarizes the semantic information contained in citations into relations within the graph. This organization effectively reflects the relationships between two academic papers and the advancement of the AI research field. Such organization aids LLMs in capturing the current progress of research, thereby enhancing their creativity. Experimental results demonstrate the effectiveness of our approach in generating novel, clear, and effective research ideas.
1 Introduction
The generation of creative ideas is a vital aspect of scientific research, fueling technological innovation and breakthroughs. Traditionally, this process has been human-driven, requiring researchers to consult extensive literature, summarize research trends, and propose improvements or optimizations as new research directions. However, in the rapidly evolving field of artificial intelligence, the increasing volume of literature and the complexity of interdisciplinary research present challenges. Researchers are constrained by the vastness of scientific literature and intricate citation relationships, demanding significant time and effort. Consequently, the overwhelming number of related papers makes this task particularly difficult, highlighting the urgent need for efficient idea generation and innovation.
Recent advancements in large language models (LLMs), such as ChatGPT [^1] https://chatgpt.com/, Claude [^2] https://claude.ai/, and LLaMA Touvron et al. (2023a, b), have demonstrated substantial potential in fostering innovation. Existing studies have shown that leveraging the specialized knowledge and exceptional natural language generation abilities acquired through large-scale pretraining, LLMs are capable not only of understanding and generating complex academic content Si et al. (2024); Kumar et al. (2024) but also excelling in coordinating constructing implicit chains of thought and revealing subtle connections between papers Wang et al. (2024b); Li et al. (2024). Utilizing LLMs to assist researchers in generating new ideas is of significant importance for enhancing research productivity.
However, current methods of idea generation based on LLMs still face three main challenges: 1) When LLMs are tasked with generating ideas, they are provided with a substantial amount of research literature and asked to summarize development trends based on the sequence and citation relationships between papers. However, the semantic information embedded within these citation relationships is not adequately considered. For example, a cited paper could be used as a baseline or as a counterexample, which makes it easy for LLMs to be influenced by less relevant works, potentially leading to ideas that lack logical coherence and technical innovation. 2) The approach of prompting LLMs to analyze citation relationships and establish chains of thought only summarizes development trends in a simple linear structure. This one-dimensional data structure oversimplifies the summary of research trends. In fact, the relationships between academic papers can be represented as a complex graph structure through knowledge graphs. Summarizing and mining more intricate relationships can yield more diverse research ideas. 3) Existing methods for generating research ideas using LLMs typically rely on human evaluation and prompt-based assessments, lacking an objective evaluation mechanism for assessing, providing feedback, and refining the generated ideas.
To address the aforementioned challenges and limitations, this paper proposes a method for generating novel AI research ideas based on knowledge graphs and large language models, specifically aimed at the rapidly evolving field of artificial intelligence, which is characterized by a large volume of open-access papers. The proposed method, Graph of AI Ideas (GoAI), works as follows. First, GoAI searches for relevant literature based on the topic under investigation and selects the most pertinent papers as key references. Once the key references are identified, GoAI analyzes their cited references and the articles that cite them, constructing a knowledge graph, the GoAI Graph, that represents the research trends of the topic. In this knowledge graph, entities represent papers, and the relations between entities reflect the semantic descriptions of citation relationships within the papers, such as “based on…”, “support by…”, “in contrast …”, etc. Furthermore, these relationships also incorporate the position where the references are cited in the original text to assess the importance of different citing papers. After constructing the graph, we employ an agent based on large language models, the GoAI Agent, to collaborate with the knowledge graph. Using a beam search algorithm, the large language model dynamically explores the graph, investigating multiple meaningful development paths that may lead to new research ideas. These paths are explored based on the entities, the semantic relationships, and the importance of these relationships within the graph. The model then generates new research ideas, performs novelty checks, and iteratively refines them. Further evaluation confirms the effectiveness of our method.
The main contributions of this paper are as follows:
  1. 1.
    We propose a novel academic knowledge graph, GoAI, and its construction method, using the AI field as an example. This graph not only represents citation relationships between papers but also captures the semantic information contained within these citations and differentiates the importance of various references.
  2. 2.
    We introduce the GoAI Agent, a research idea generation agent that collaborates with the knowledge graph. By enabling the large language model to dynamically explore the AI idea graph, the agent analyzes trends and generates new ideas based on this exploration.
  3. 3.
    We present an evaluation method for assessing the research ideas generated by large language models for assessing, providing feedback, and refining. Experimental results demonstrate the effectiveness of the GoAI method we propose.
2 Related Works
2.1 AI-based Scientific Discovery
Artificial intelligence has become a powerful tool for advancing research processes, reshaping scientific discovery by reducing the time needed for experimental iterations, and assisting researchers in various capacities Yu et al. (2024); Abramson et al. (2024); de Haan et al. (2024); Tan et al. (2024). The development of large language models, with their advanced natural language processing capabilities and ability to integrate interdisciplinary information, enables more efficient literature analysis and the generation of novel hypotheses Kumar et al. (2024); Si et al. (2024). Prior research has enhanced novelty by iteratively comparing AI-generated ideas with existing literature Wang et al. (2024a), while other studies have focused on linking academic knowledge through core papers, expanding understanding via academic graphs and entity-centered knowledge stores Baek et al. (2024); Wang et al. (2024b); Li et al. (2024). Some studies use multi-agent systems based on LLMs to foster the generation of research ideas in autonomous scientific discovery Su et al. (2024). In this paper, we adopt a paradigm for generating research ideas based on academic graphs derived from core papers. Unlike previous works, we represent citation relationships between academic papers and summarize development trends through a knowledge graph, rather than through research trend chains. Furthermore, we introduce the concept of constructing a knowledge graph that incorporates semantic information and citation importance, providing richer information for analyzing relationships between papers.
2.2 Graph-augmented LLMs
The integration of large language models with knowledge graphs, which store information in natural language, is an emerging research area aimed at fully utilizing the explicit and updatable external knowledge within structured graphs to enhance the capabilities of large language models Pan et al. (2024); Huang et al. (2024), particularly in tasks such as reasoning Luo et al. (2024b, c) and rule completion Luo et al. (2024a) and benchmarking Ruan et al. (2025); Zhu et al. (2024); Ruan et al. (2024b, a) . To input the structured knowledge from knowledge graphs into large language models, some methods transcribe entity-relation-entity triples from the knowledge graph into natural language text that is understandable by the language model, which is then provided as a prompt to the LLM Sun et al. (2024), similar to retrieval-augmented generation. Other methods directly input graph information processed by approaches such as Graph Neural Networks (GNNs) into the large language model, utilizing the implicit graph information for enhancement Huang et al. (2024); Mavromatis and Karypis (2024). In this paper, since the literature used to construct the graph is already in natural language and the relations within the graph contain semantic information in natural language, we adopt the method of transcribing the graph into textual prompts for input into the large language model to achieve the collaboration between the graph and the model.
3 Methodology
Figure 1: The framework of GoAI. The framework consists of four stages: (1) Literature Search and Filtering, (2) GoAI Graph Construction, (3) Idea Generation, and (4) Novelty Evaluation.
3.1 GoAI Framework
In this section, we will provide a detailed description of the GoAI method framework. As shown in Figure 1, the framework consists of four stages: (1) Literature Search and Filtering, (2) GoAI Graph Construction, (3) Idea Generation, and (4) Novelty Evaluation.
First, given a research topic, GoAI searches the online literature database for relevant papers and identifies key reference papers based on their ranking. Based on the citation relationships of these key reference papers, the graph is extended both forward and backward to reflect various research trends in the field. Then, using LLM-as-agent, the knowledge graph is analyzed to predict future research directions. The generated ideas undergo novelty checks via structured-thinking evaluation, with iterative refinement and integration leading to the final ideas. The prompts used in this section can be found in our supplementary materials.
3.2 Literature Search and Filtering
We use the Semantic Scholar API [^3] https://www.semanticscholar.org/product/api to search for papers in the online paper database. Given an initial research topic, we retrieve the top-ranked papers, which we refer to as key references, forming the foundation for building the graph. Starting from the corresponding key reference papers, GoAI constructs the graph by extending in both forward and backward directions based on citation relationships. The backward extension is derived from the reference list of the key reference, tracing the origins of the paper’s ideas. The forward extension is achieved by exploring which papers in the literature cite the key reference paper, representing the subsequent development of the idea.
For the papers in the backward extension, we rank them based on their cosine similarity to the initial research topic and the abstract of the key reference paper. The top $K$ papers are selected and added to the graph, representing one step of the process. For each new paper added to the graph, the same rule is applied, extending backward based on its reference list until the number of backward steps reaches the set upper limit $N$. A similar procedure is applied in the forward direction, extending $N$ steps. If no references relevant to the original research topic are found during the extension process, GoAI terminates the search. Under ideal conditions, the number of papers retrieved is $O(2K^{N})$.
Figure 2: An example workflow of graph exploration and idea generation. The paper “Tree of Thoughts” is the key reference. A more detailed sub-graph example can be found in our supplementary materials.
3.3 GoAI Graph Construction
3.3.1 GoAI Graph Overview
The GoAI graph is used to represent the relationships between papers that have citation relationships, including the citing and cited papers, the importance of the citation (measured by the position where the reference appears in the original text), and the semantic information contained in the citation. The general representation of the GoAI graph is a structured <paper, citation position, citation semantics, paper> quadruple. The citation position and citation semantics together form the relation field in traditional knowledge graphs, which is used for subsequent analysis of the graph. To construct the GoAI graph, it is necessary to separately obtain the papers as entities and the citation positions and citation semantics as relations.
3.3.2 Relation Definition and Extraction
After determining the entities to be included in the graph, we proceed with constructing the relations between these entities. For the relationships between different papers, we define the citation relationships between papers as one of the following five semantic categories:
  • Based on and Extension (B&E): The paper is based on, extends, applies, or is inspired by the cited paper, or generalizes its theories and methods.
  • Support and Supplement (S&S): The paper supports the cited paper’s work through citation, reuse, supplementation, or indirect connection.
  • Contrast and Alternative (C&A): The paper compares itself with the cited paper, proposes alternatives, or summarizes its content from a different perspective.
  • Question and Refutation (Q&A): The paper questions, corrects, or refutes the content of the cited paper.
  • Simple Mention or Irrelevant (M/I): The paper merely mentions the cited paper, or there is little to no direct relevance between the two.
We use the SciPDF Parser [^4] https://github.com/titipata/scipdf_parser to parse the paragraphs from the paper documents, prompting the LLM to extract the cited references from the original paragraphs and determine which of the five semantic relationships best describes the citation.
3.4 Idea Generation
The process of generating ideas from the graph is divided into two steps: graph exploration and idea generation. Figure 2 illustrates an example workflow of graph exploration and idea generation.
3.4.1 Graph Exploration
In the graph, the entities are papers, and the relations are represented as pairs of citation positions and citation semantics. At the beginning of the $D$-th iteration, each path consists of $12D-$ 1 quadruples, i.e., $p=\{(e^{d}_{s},r^{d},e^{d}_{o})\}^{2D-1}_{d=1}$, where $e_{s}$ and $e_{o}$ represent the subject and object entities, respectively, and $r$ denotes the specific relation between them. The pairs $(e^{d}_{s},r^{d},e^{d}_{o})$ and $(e^{d+1}_{s},r^{d+1},e^{d+1}_{o})$ are interconnected. The sets of endpoint entities and relations of all paths in the path set $P$ are denoted as $E^{D-1}=\{e^{d-1}_{1},e^{d-1}_{2},\cdots,e^{d-1}_{N}\}$ and $R^{D-1}=\{r^{d-1}_{1},r^{d-1}_{2},\cdots,r^{d-1}_{N}\}$ , respectively.
The exploration phase of the $D$-th iteration aims to utilize the LLM to identify the top-N most relevant entities $E^{D}$ from the neighboring entities of the current top-N endpoint entity set $E^{D-1}$, and use $E^{D}$ to extend the top-N paths in the path set $P$. The endpoint entities here can be either the head or tail entities of the path, corresponding to backward and forward extensions, respectively. To address the complexity of processing a large number of neighboring entities using the LLM, inspired by the approach of ToG Sun et al. (2024), this paper implements a two-step exploration strategy: first, explore the important relations, and then use the selected relations to guide entity exploration.
Relation Exploration The relation exploration is a beam search process with a depth of 1 and a width of $N$, from $E^{D-1}$ to $R^{D}$ The entire process can be broken down into two steps: search and pruning. We use the GoAI Agent, an LLM-based agent, to automatically complete this process.
  • Search Step At the beginning of the $D$-th iteration, the relation exploration phase first associates each path $p_{n}$ with the candidate relations $R^{D}_{cand,n}$ that are related to the endpoint entity $e_{n}^{D-1}$. These relations are then aggregated into $R^{D}_{cand}$. In the case of Figure 2, $E^{1}=$Tree of Thoughts} and $R^{1}$ denotes the set of all relations linked to Tree of Thoughts inwards or outwards.
  • Pruning Step Once the candidate relation set $R^{D}_{cand}$ and the extended candidate paths $P_{cand}$ are obtained from the relation search, the LLM-based GoAI Agent can be used to select the top-N new paths $P$ that have the relation $R^{D}$ as the endpoint from $P_{cand}$, based on the textual information of the query and the candidate relations $R^{D}_{cand}$. The pruning process can be performed by prompting the LLM, allowing the GoAI Agent to adapt well to different graphs without any training cost. As shown in Figure 2, the LLM selects relations (Background, B&E), (Introduction, C&A), (Introduction, B&E) out from all relations linked to the entity Tree of Thoughts in the first iteration. Since Tree of Thoughts is the only topic entity, the candidate paths are updated as {(Tree of Thoughts, (Background, B&E)), (Tree of Thoughts, (Tree of Thoughts, C&A)),(Tree of Thoughts, (Introduction, C&A))}.
Entity Exploration Similar to relation exploration, entity exploration is also a beam search process performed by the LLM from $R_{D}$ to $E_{D}$, consisting of two steps: Search and Prune.
[…]
  • Pruning Step Since each entity in the candidate set $E^{D}_{cand}$ is expressed in natural language, we can use the GoAI Agent to select the top-N new paths from $P_{cand}$, where these paths end with the entity $E^{D}$. As shown in Figure 2, the current paths $P$ are updated as {(Tree of Thoughts, (Background, B&E), Self-Consistency), (Tree of Thoughts, (Introduction, C&A), CoT),(Tree of Thoughts, (Introduction, C&A), CPO),(Tree of Thoughts, (Introduction, B&E), Diagram of Thought),(Tree of Thoughts, (Introduction, B&E), ControlLLM)}.
After performing the above two explorations, we construct the top-N new paths $P=\{p_{n}\}_{n=1}^{N}$, where the length of each path is increased by 2.
3.4.2 Idea Generation with paths
After obtaining the current path $p$ through the exploration process, we prompt the LLM to evaluate the development trends of the papers in the current path, which guides the generation of new ideas. For each generated path, the model predicts the potential future trends. Once the future trends are identified, we continue to prompt the LLM to elaborate on the motivations, novelty, and methodology, ultimately solidifying the idea. Through this gradual process, more detailed ideas can be generated. We also use the novelty checking agent introduced in Section 3.5 to assess candidate ideas and iteratively refine them.
3.5 Novelty Evaluation with GoAI-CoT-Reviewer
Recent research indicates that a significant factor influencing the capabilities of LLMs is the systematicity and structured nature of the reasoning process. Systematicity refers to the model’s ability to perform multi-stage reasoning rather than arriving at a conclusion in a single step. Structuring refers to the model’s ability to clearly identify its current stage of thinking and understand the primary task to be addressed in each stage. OpenAI’s o1 Zhong et al. (2024) effectively addressed these issues by enabling the model to independently perform systematic and structured reasoning through language, while LLaVA-CoT Xu et al. (2025) enhanced the model’s step-by-step reasoning ability via supervised fine-tuning.
Given that the evaluation of paper idea novelty requires the model to deeply analyze the content of the idea and provide a rating, the model needs to engage in profound, step-by-step thinking. Therefore, we evaluate the novelty of generated ideas through structured, step-by-step assessment. By decomposing the evaluation of paper novelty into three structured stages, we constructed an evaluation model, GoAI-CoT-Reviewer, to evaluate the novelty of ideas for assessing, providing feedback, and refining.
3.5.1 Step-by-Step Evaluation Stages
Figure 3: Processing flow of novelty evaluation for GoAI-CoT-Reviewer.
As shown in Figure 3, inspired by the approach of LLaVA-CoT Zhong et al. (2024) and incorporating the thought process of human reviewers, we decompose the evaluation of paper novelty into three structured stages: summary, analysis, and scoring.
Summary Stage: In this initial stage, GoAI-CoT-Reviewer is provided with an idea and a corresponding paper abstract based on this idea. GoAI-CoT-Reviewer briefly summarizes the main content and methods expressed by the idea based on this information.
Analysis Stage: Based on the summarized content and methods, GoAI-CoT-Reviewer analyzes the strengths and weaknesses of the idea, which will serve as the basis for the subsequent scoring.
Scoring Stage: Based on the results obtained from the first two stages, GoAI-CoT-Reviewer provides a numerical score to assess the novelty of the idea.
GoAI-CoT-Reviewer uses dedicated tags (e.g., <Summary>...</Summary>) to mark each stage, indicating the beginning and end of each phase. These tags help the model maintain clarity throughout the reasoning process. Unlike traditional CoT Wei et al. (2023) reasoning, which allows the model to freely think, our approach promotes structured thinking by first summarizing the input information, then performing a detailed analysis, and finally arriving at a score. To achieve this, we construct a training dataset by progressively generating responses using open review comments from the Openreview platform and then train the evaluation model using supervised fine-tuning. After training, the model autonomously selects these labels as needed, completing all stages within a single process. This structured approach enables the model to independently manage its reasoning process, improving its performance on complex tasks that require multi-step analysis.
3.5.2 Data Preparation and Model Training
We use the Openreview API v2555https://docs.openreview.net/reference/api-v2 to scrape open-source AI conference papers and corresponding review comments submitted on the Openreview platform, including both accepted and rejected papers. To ensure the quality of the papers and review comments to some extent, we scraped papers and reviews from top AI conferences such as ICLR 2022, NeurIPS 2022, and ICLR 2023, with a total of over 35k samples in the training dataset.
Following the approach of the LLAVA-CoT-100K dataset Xu et al. (2025), we use the summary_of_the_paper field from the review comments as the golden answer for the Summary stage, the summary_of_the_paper and strengths_and_weaknesses fields as the golden answers for the Analysis stage, and the technical_novelty_and_significance field as the golden answer for the Scoring stage to construct a supervised fine-tuning dataset. We fine-tune the base model using LoRA to obtain the GoAI-CoT-Reviewer model, which is used for novelty evaluation.
4 Experiments
4.1 Setup
Due to the large number of papers and their lengthy nature, and to reduce costs, we use the GLM-4-Flash API [^6] https://bigmodel.cn/dev/activities/free/glm-4-flash as the backend large language model. The GoAI-CoT-Reviewer evaluation model is built on the Llama 3.1-8B-Instruct model Dubey et al. (2024) and fine-tuned through LoRA. Training is conducted using NVIDIA A100 GPUs, with a LoRA rank set to 8, an input truncation length of 2048, and the number of epochs set to 3.
We use Idea Arena Li et al. (2024), a pairwise evaluation system, to assess the novelty of ideas generated by large language models. Idea Arena utilizes a round-robin competition to calculate the Elo score for each idea generation method. For a given topic, we ask large language model judges to rank the generated ideas based on dimensions such as novelty, significance, clarity, and expected effectiveness. The Elo scores for each method are then calculated using the win-loss-tie records obtained from these evaluations Zheng et al. (2023); Zhao et al. (2024). A detailed description of the evaluation metrics can be found in our supplementary materials.
Method Novelty Significance Effectiveness Clarity Feasibility Average
GoAI (ours) 3.83 3.27 3.6 3.02 3.22 3.388
Chain of Ideas Li et al. (2024) 3.21 3.17 3.77 2.95 3.24 3.268
AI-Scientist Lu et al. (2024) 2.79 2.88 2.58 3.05 2.83 2.826
ResearchAgent Baek et al. (2024) 2.73 2.87 2.67 2.94 2.98 2.838
Vanilla LLM 2.44 2.81 2.38 3.04 2.85 2.704
Table 1: Evaluation results of idea generation.
Method Pearson Corr.
Direct Generation 0.0058
Multi-turn Dialogue 0.3119
Structured Thinking (ours) 0.5271
Table 2: The impact of SFT dataset on evaluation model.
Figure 4: Evaluation results of idea generation.
Figure 5: Case study for the idea generation result of our GoAI compared with Chain of Ideas Li et al. (2024) in the same topic “Eliminate Hallucinations in Large Language Models”. The paper number from 1 to 6 represent Qi et al. (2021), Min et al. (2020), Zhou et al. (2021), Ahn et al. (2022), Izacard et al. (2022), and Chang et al. (2024), respectively.
4.2 Main Results
4.2.1 Human-Model Agreements of Evaluation
Table 2 presents the Pearson correlation coefficient between our evaluation model and human evaluations. The results show that our evaluation model maintains a good positive correlation with human reviewers’ assessments, making it a valuable tool for providing feedback in the idea generation process.
4.2.2 Novelty of Generated Ideas
Figure 4 and Table 1 display a comparison of the creative idea generation results between our proposed method and other idea generation approaches. Overall, compared to all other automated methods, our GoAI demonstrates superior performance across evaluations. Notably, GoAI excels in generating ideas with novelty and significance, highlighting its exceptional capability in creative idea generation. In other areas, considering that no targeted optimization was performed, the gap between GoAI and other methods is expected. Nevertheless, even in these aspects, GoAI still exhibits strong performance, allowing it to outperform other methods in terms of average scores.
4.3 Ablation Study for Evaluation Model
To investigate the impact of the structured thinking process SFT training dataset on the evaluation model, we compared it with models trained using raw data directly and using datasets in a multi-turn dialogue format. The evaluation metric used is the Pearson correlation coefficient with human ratings. The dataset trained with raw data takes the form of a QA question-answering format, where the input is the title and abstract of a paper, and the output is the novelty score of the paper. The multi-turn dialogue format dataset consists of several rounds of QA about the paper’s summary, strengths and weaknesses, and score. Table 2 presents the Pearson correlation coefficient results for the models trained with different datasets. The results show that using the structured CoT thinking process significantly improves the correlation between the evaluation model and human assessments, consistent with the findings in Xu et al. (2025).
4.4 Case Study
Figure 5 presents a case study comparing our proposed GoAI method with the Chain of Ideas Li et al. (2024). The topic used for generating ideas is “Eliminate Hallucinations in Large Language Models,” and the starting point for constructing path analysis of research trends is the same paper, “Retrieval Augmentation Reduces Hallucination in Conversation.” From the paper paths, it can be observed that our method, by analyzing the semantic information of citations, is able to retrieve papers more closely related to the topic. The generated idea also aligns more closely with the paper path, indicating that the idea generation comes from a thorough analysis of the development trends in the field. In contrast, the paths constructed by the Chain of Ideas method include a paper that is unrelated to the topic being studied, titled “cTBLS: Augmenting Large Language Models with Conversational Tables.” The generated idea is also unrelated to the papers in the path, indicating that the papers retrieved by this method are easily influenced by irrelevant citations, and the generated idea may originate from the model’s own pre-trained knowledge or hallucinations.
5 Conclusion
In this paper, we introduced the Graph of AI Ideas (GoAI), a framework designed to enhance the capability of LLMs to generate research ideas in the field of AI. By organizing ideas from existing papers into a structured graph, GoAI effectively captures the progressive relationships and incremental developments between studies. The use of an LLM-based agent for path searching in the graph helps LLMs to understand and analyze current research progress, uncover connections between existing studies, and thereby enhance their creative potential.
To evaluate and refine the automatically generated ideas, we also proposed GoAI-CoT-Reviewer, an evaluation model with structured thinking capabilities. This model evaluates the novelty of ideas through a reasoning process similar to that of human reviewers, providing feedback highly consistent with human assessments. Experimental results demonstrate that GoAI consistently outperforms other methods and is capable of generating ideas comparable to human creativity.
Ethical Statement
This study constructs a knowledge graph using publicly available research paper data to generate novel ideas. All data were sourced from open academic repositories, with no personal or sensitive information involved. Data processing followed the respective terms of use, and no privacy concerns arose. The authors declare no conflicts of interest, and all analyses were conducted transparently using automated methods. The study complies with ethical guidelines for responsible data use in academic research.